Help me solve: 12T+7=6+11
I've come up with 6t +7=11
6t/6=4/6 cant figure out from here

Answer by edjones(8007)

12t+7=6+11
12t+7=17
12t=10 subtract 7 from each side.
t=10/12 divide each side by 12
t=5/6 simplify
.
Ed